NG 2019> Assessment 2: Stem & Leaf, Histograms, Percentiles Due in 13 Ages Number of students 15-181 。 19-22 23-26 27-30 10 35-38 Find the relative frequency for the class with lower class limit 31 Relative Frequency Give your answer as a percent, rounded to two decimal places ersion Points possible: 1 This is attempt 1 of 3. Submit t/showtest.phplactionzskip&to 8
Answer #1

Total of the frequencies is

6+3+2+10+4+7 = 32

And total frequencies for class limit lower than 31 is:

6+3+2+10 = 21

So relative frequency becomes (21÷32)*100 = 65.63